OpenStudy (anonymous):

let g be the function given by g(x)=intergral from 0 to x of sin(t^2)dt for -1

OpenStudy (amriju):

find its derivative...then find where its negative..its decreasing there...for differentiation use newton leibnitz rule..:variable limits form http://en.wikipedia.org/wiki/Leibniz_integral_rule

OpenStudy (anonymous):

hey thanks man. i really appreciate it

OpenStudy (anonymous):

Also note that \[\int_0^x \sin (t^2)dt=F(x)-F(0) \] so if you need to differentiate it with respect to \(x\) you will be left with \(f(x)=\sin(x^2)\) =g'(x)

OpenStudy (amriju):

u're welcome..:)

OpenStudy (anonymous):

ahh. good point spacelimbus.